# Tiny Time Mixers (TTMs): Fast Pre-trained Models for Enhanced Zero/Few-Shot Forecasting of Multivariate Time Series

### 17 Jan 2024 v3

@@ 1 Introduction
Multivariate time series (TS) forecasting entails predicting future values for multiple interrelated time series based on their historical data. This field has advanced significantly, applying statistical and machine learning (ML) methods Hyndman and Athanasopoulos (2021) across domains like weather, traffic, retail, and energy. In general, each time series represents a variable or channel111“Channel” refers to the individual time series in multivariate data (i.e., a multivariate TS is a multi-channel signal).. In certain applications, non-forecasting variables, categorized as controllable and uncontrollable external factors, impact the variables to forecast. We term these non-forecasting variables as exogenous, and the variables requiring forecast as target variables.
Related Work: Recent advances in multivariate forecasting have been marked by the advent of transformer-based Vaswani et al. (2017) approaches, exemplified by models like PatchTST Nie et al. (2023), Autoformer Wu et al. (2021), Informer Zhou et al. (2021), and FEDFormer Zhou et al. (2022). These models have demonstrated notable improvements over traditional statistical and ML methods. Furthermore, architectures based on MLP-Mixer Tolstikhin et al. (2021), such as TSMixer Ekambaram et al. (2023), have emerged as efficient transformer alternatives, boasting 2-3X reduced compute and memory requirements with no accuracy compromise compared to their transformer counterparts. However, none of these advanced approaches have successfully demonstrated the ability to create general pre-trained models that can successfully transfer the learning to unseen target TS dataset, in a similar way as popularly witnessed in NLP and vision tasks. This is very challenging in the TS domain due to the diverse nature of the datasets across applications and the limited public availability of TS data for pre-training. There are existing self-supervised pre-training TS approaches using masked modeling and contrastive learning techniques such as SimMTM Dong et al. (2023) and TF-C Zhang et al. (2022) that offer transfer learning between two datasets when carefully selected based on the dataset properties. However, they fail to provide universal transfer learning capabilities across datasets. Consequently, there has been a recent growing trend to employ pre-trained large language models (LLMs) for TS forecasting, treating it as a cross-domain transfer learning task. These universal cross-transfer approaches, specifically recent works such as LLMTime Gruver et al. (2023) and GPT4TS Zhou et al. (2023) yield promising results in few/zero-shot forecasting approaches. These models are bootstrapped from GPT-2/3 or LLAMA-2 with suitable tokenization strategies to adapt to time-series domains.
@@
In particular, TTM is pre-trained using multiple public datasets ($\sim$ 244M samples) from the Monash data repository [^2] Accessible at https://forecastingdata.org/ Godahewa et al. (2021)). Note that the datasets exhibit considerable diversity in terms of characteristics, such as the different domains, temporal resolution [^3] Resolution refers to the sampling rate of the input time series (e.g., hourly, 10 minutes, 15 minutes, etc.) (spanning from second to daily), lengths, and number of channels. Pre-training on such heterogeneous datasets cannot be handled directly by TSMixer or existing state-of-the-art (SOTA) models. Hence, TTM proposes the following enhancements to the TSMixer architecture: (i) Adaptive Patching across layers, considering the varied suitability of patch lengths for different datasets, (ii) Dataset Augmentation via Downsampling to increase coverage and samples across different resolutions, (iii) Resolution Prefix Tuning to explicitly embed resolution information in the first patch, facilitating resolution-conditioned modeling, particularly beneficial in scenarios with short history lengths. Moreover, our approach leverages multi-level modeling, where TTMs are first pre-trained in a channel-independent way and then seamlessly integrate channel mixing during fine-tuning to model target data-specific channel-correlations and exogenous infusion
@@ 2.1 Multi-level Modeling
TTM follows a multi-level architecture consisting of four key components (see Figure 1(a)): (1) The TTM Backbone is assembled using building blocks derived from the efficient TSMixer architecture Ekambaram et al. (2023). TSMixer is based on simple MLP blocks that enable mixing of features within patches, across patches and channels, surpassing existing transformer-based TS approaches with minimal computational requirements. Since TSMixer is not targeted to handle multi-resolution data, we introduce various novel enhancements to it as explained later. (2) TTM Decoderfollows the same backbone architecture but is considerably smaller in size, approximately 10-20% of the size of the backbone, (3) Forecast Headconsists of a linear head designed to produce the forecast output, and (4) Optional Exogenous Mixer serves to fuse exogenous data into the model’s forecasting process. This multi-level model refactoring is required to dynamically change the working behavior of various components based on the workflow type, as explained in Section 3. In addition to the above primary components, we also have a preprocessing component as explained next.
@@ 2.2 Pre-processing
As shown in Figure 1(a) with colorless blocks, the historical time series $\bm{X}$ is first normalized per instance to have zero mean and unit standard deviation for each channel dimension, to tackle any possible distribution shifts Nie et al. (2023); Ekambaram et al. (2023). This process is reversed at the end before computing the loss. The normalized data $\overline{\bm{X}}$ is subsequently patched $\bm{X}_{p}\in\mathbb{R}^{c\times n\times pl}$ into $n$ non-overlapping windows, each of length $pl$ and then, passed to the TTM backbone. Patching, as introduced in Nie et al. (2023), has proven to be highly valuable for forecasting. Its effectiveness lies in preserving local semantic information, accommodating longer history, and reducing computation.
@@ 3.1 Pre-training Workflow
In the pre-training stage, we train the model on a large collection of public datasets from the Monash data repositoryGodahewa et al. (2021). Since the primary focus of TTM is forecasting, pre-training is modeled with a direct forecasting objective. TTM is first pre-trained in a univariate fashion with independent channels on all the existing datasets. Due to varied channel counts in pre-training datasets, modeling multivariate correlations is not feasible here; it is addressed later during fine-tuning. Multivariate pre-training datasets are initially transformed into independent univariate time series $(\bm{X}_{1},\cdots,\bm{X}_{N})\in\mathbb{R}^{c(=1)\times sl}$. These are pre-processed (Section 2.2), and subsequently fed into the TTM backbone for multi-resolution pre-training. The output of the backbone $\bm{X}_{h}^{L}\in\mathbb{R}^{(c=1)\times n\times hf}$ is passed through the decoder and forecast head to produce the forecast $\hat{\bm{Y}}\in\mathbb{R}^{(c=1)\times fl}$ which is then reverse-normalized to bring back to the original scale. We pre-train the TTM with mean squared error (MSE) loss function calculated over the forecast horizon: $\mathcal{L}=||\bm{Y}-\hat{\bm{Y}}||_{2}^{2}$. Thus for a given input context length $sl$ and forecast length $fl$, we get a pre-trained model capturing the common temporal forecasting dynamics and seasonal patterns as observed in the overall pre-training data.
3.1.1 Multi-Resolution Pre-training via TTM Backbone
@@
The TTM backbone consists of $L$ levels, each comprising $M$ TTM blocks with identical patch configurations. A TTM block in the $i$-th level, $i=1,\ldots,L$, receives the processed data $\bm{X}_{h}^{(i-1)}\in\mathbb{R}^{c\times n\times hf}$ from the earlier block. Each TTM block is further comprised of a patch partition model, a vanilla TSMixer block, and a patch merging block. Patch Partition block at every level $i$ increases the number of patches by a factor of $K_{i}$ and reduces the patch dimension size by the same factor by reshaping $\bm{X}_{h}^{(i-1)}\in\mathbb{R}^{c\times n\times hf}\text{to }\bm{X}_{h}^{i}\in\mathbb{R}^{c\times(n\cdot K_{i})\times(hf/K_{i})}$, where $K_{i}=2^{(L-i)}$. Note that, we set $hf=2^{m}$ for some integer $m\geq L$. Then, TSMixer is applied to the adapted data $\bm{X}_{h}^{i}$. Finally, the output from TSMixer is again reshaped to its original shape (i.e., $\mathbb{R}^{c\times n\times hf}$) in the patch merging block. Note that, as we go deeper into the network, the number of patches decreases while the patch dimension size increases leading to adaptive patching which helps in better generalization as we pre-train with multiple datasets together. This idea of adaptive patching is popular and very successful in the vision domain (E.g. Swin transformers Liu et al. (2021) and we are the first to port it successfully to the time-series domain to resolve multi-resolution issues in pre-training with diverse TS datasets. Figure 1(b) shows the TTM backbone for $L=3$ and $M=2$. Please note that adaptive patching is enabled only in the backbone and not in the decoder, which is designed to be very lightweight.
@@
Resolution Prefix Tuning: This technique explicitly learns and incorporates a new patch embedding as a prefix into the input data based on the input resolution type (see Figure 1(b)). Similar to the concept of prefix tuning Li and Liang (2021), this approach provides an explicit signal to the model about the resolution type for resolution-conditioned modeling. First, we map every resolution to a unique integer, which is then passed through an embedding layer to project it to the hidden dimension, $hf$. Subsequently, we expand the embedding across all channels to have a representation of shape $c\times 1\times hf$. This module is optional for the TTM backbone, particularly beneficial when the context length ($sl$) is short. In these scenarios, automatically detecting the resolution becomes a challenge for the model. Hence, by explicitly fusing the resolution information as a prefix, we can enhance the model’s ability to learn effectively across resolutions.
@@ 3.2 Fine-tuning Workflow
Zero-shot Few-shot (5%) $fl$ TTM TTM GPT4TS PatchTST TSMixer TimesNet Dlinear FEDFormer Autoformer ETTH1 96 0.365 0.366 0.543 0.557 0.554 0.892 0.547 0.593 0.681 192 0.393 0.391 0.748 0.711 0.673 0.940 0.720 0.652 0.725 336 0.415 0.421 0.754 0.816 0.678 0.945 0.984 0.731 0.761 720 0.538 - - - - - - - - ETTH2 96 0.285 0.282 0.376 0.401 0.348 0.409 0.442 0.390 0.428 192 0.341 0.338 0.418 0.452 0.419 0.483 0.617 0.457 0.496 336 0.383 0.383 0.408 0.464 0.389 0.499 1.424 0.477 0.486 720 0.441 - - - - - - - - ETTM1 96 0.413 0.359 0.386 0.399 0.361 0.606 0.332 0.628 0.726 192 0.476 0.402 0.440 0.441 0.411 0.681 0.358 0.666 0.750 336 0.553 0.424 0.485 0.499 0.467 0.786 0.402 0.807 0.851 720 0.737 0.575 0.577 0.767 0.677 0.796 0.511 0.822 0.857 ETTM2 96 0.187 0.174 0.199 0.206 0.200 0.220 0.236 0.229 0.232 192 0.261 0.240 0.256 0.264 0.265 0.311 0.306 0.394 0.291 336 0.323 0.299 0.318 0.334 0.314 0.338 0.380 0.378 0.478 720 0.436 0.407 0.46 0.454 0.410 0.509 0.674 0.523 0.553 Weather 96 0.154 0.152 0.175 0.171 0.188 0.207 0.184 0.229 0.227 192 0.203 0.198 0.227 0.230 0.234 0.272 0.228 0.265 0.278 336 0.256 0.250 0.286 0.294 0.287 0.313 0.279 0.353 0.351 720 0.329 0.326 0.366 0.384 0.365 0.400 0.364 0.391 0.387 Electricity 96 0.169 0.142 0.143 0.145 0.147 0.315 0.150 0.235 0.297 192 0.196 0.162 0.159 0.163 0.172 0.318 0.163 0.247 0.308 336 0.209 0.184 0.179 0.175 0.190 0.340 0.175 0.267 0.354 720 0.264 0.228 0.233 0.219 0.280 0.635 0.219 0.318 0.426 Traffic 96 0.518 0.401 0.419 0.404 0.408 0.854 0.427 0.670 0.795 192 0.548 0.425 0.434 0.412 0.421 0.894 0.447 0.653 0.837 336 0.55 0.437 0.449 0.439 0.477 0.853 0.478 0.707 0.867 720 0.605 - - - - - - - - TTM Zero-shot % IMP 1% 3% 2% 31% 6% 25% 32% TTM 5% Few-shot % IMP 12% 14% 12% 38% 17% 32% 38% Table 1: Zero-shot and Few-shot 5% performance (MSE) of TTM and all SOTA models on D1 datasets for varying forecast lengths ($fl$). Bold and underscore denote the best and the second best. For longer $fl$ (720), some datasets fail to generate fine-tuning samples with just 5% of the data, denoted as ‘-’. The last two rows show relative improvement (IMP) of TTM zero-shot and few-shot w.r.t. SOTA few-shot results.
@@ 4.1 Experimental Setting
Datasets: Pre-training employs a subset of the Monash data hub Godahewa et al. (2021) of size $\sim 244$M samples. We specifically exclude datasets (like yearly, monthly) as they do not possess sufficient length for the long-term forecasting task. Moreover, we remove all the datasets that we utilize for evaluation (i.e., weather, electricity, and traffic). For zero/few-shot evaluation we consider seven public datasets (D1): ETTH1, ETTH2, ETTM1, ETTM2, Weather, Electricity & Traffic as popularly used in most prior SOTA works Zhou et al. (2021); Nie et al. (2023). Since these datasets do not contain any exogenous variables nor exhibit cross-channel correlation benefits, we incorporate four other datasets (D2) for separately validating the efficacy of the decoder channel mixing and exogenous mixer module: bike sharing (BS) Fanaee-T (2013), carbon capture plant (CC) Jablonka et al. (2023), and 2 more datasets from Biz-IT observability domain ITB (2023): Application (APP) and Service (SER). Refer Appendix for full data details.
SOTA Bechmarks: We benchmark TTM with the latest public SOTA forecasting models categorized as follows: [label=()] 1. LLM-based TS pre-trained models:GPT4TS Zhou et al. (2023), LLMTime Gruver et al. (2023) 2. Self-supervised pre-trained models: SimMTM Dong et al. (2023),Ti-MAE Li et al. (2023), TST Zerveas et al. (2021), LaST Wang et al. (2022), TF-C Zhang et al. (2022), CoST Woo et al. (2022) and Ts2Vec Yue et al. (2022) 3. TS transformer models:PatchTST Nie et al. (2023), FEDFormer Zhou et al. (2022), Autoformer Wu et al. (2021) 4. Other SOTA models:TSMixer Ekambaram et al. (2023), DLinear Zeng et al. (2022) and TimesNet Wu et al. (2022)
Zero-shot Few-shot (10%) Data TTM (Ours) TTM (Ours) GPT4TS PatchTST TSMixer TimesNet Dlinear FEDFormer Autoformer ETTH1 0.428 0.422 0.590 0.633 0.578 0.869 0.691 0.639 0.702 ETTH2 0.362 0.360 0.397 0.415 0.370 0.479 0.605 0.466 0.488 ETTM1 0.545 0.458 0.464 0.501 0.441 0.677 0.411 0.722 0.802 ETTM2 0.302 0.278 0.293 0.296 0.282 0.320 0.316 0.463 1.342 Electricity 0.210 0.180 0.176 0.176 0.180 0.323 0.180 0.346 0.431 Traffic 0.555 0.428 0.440 0.430 0.429 0.951 0.446 0.663 0.749 Weather 0.236 0.227 0.238 0.242 0.236 0.279 0.241 0.284 0.300 TTM Zero-shot improvements -4% -2% -7% 27% 2% 27% 39% TTM Few-shot (10%) improvements 7% 9% 4% 34% 13% 34% 45% Table 3: Zero-shot and Few-shot 10% performance (MSE) of TTM over all SOTA models on D1 datasets, averaged across 4 forecast lengths. The numbers with bold and underscore denote the best and the second best results respectively. The full table is in Appendix.

### 09 Apr 2024 v5

@@ 1 Introduction
Multivariate time series (TS) forecasting entails predicting future values for multiple interrelated time series based on their historical data. This field has advanced significantly, applying statistical and machine learning (ML) methods [?] across domains like weather, traffic, retail, and energy. In general, each time series represents a variable or channel111“Channel” refers to the individual time series in multivariate data (i.e., a multivariate TS is a multi-channel signal).. In certain applications, non-forecasting variables, categorized as controllable and uncontrollable external factors, impact the variables to forecast. We term these non-forecasting variables as exogenous, and the variables requiring forecast as target variables.
Related Work: Recent advances in multivariate forecasting have been marked by the advent of transformer-based [?] approaches, exemplified by models like PatchTST [?], Autoformer [?], Informer [?], and FEDFormer [?]. These models have demonstrated notable improvements over traditional statistical and ML methods. Furthermore, architectures based on MLP-Mixer [?], such as TSMixer [?], have emerged as efficient transformer alternatives, boasting 2-3X reduced compute and memory requirements with no accuracy compromise compared to their transformer counterparts. However, none of these advanced approaches have successfully demonstrated the ability to create general pre-trained models that can successfully transfer the learning to unseen target TS dataset, in a similar way as popularly witnessed in NLP and vision tasks. This is very challenging in the TS domain due to the diverse nature of the datasets across applications and the limited public availability of TS data for pre-training. There are existing self-supervised pre-training TS approaches using masked modeling and contrastive learning techniques such as SimMTM [?] and TF-C [?] that offer transfer learning between two datasets when carefully selected based on the dataset properties. However, they fail to provide universal transfer learning capabilities across datasets. Consequently, there has been a recent growing trend to employ pre-trained large language models (LLMs) for TS forecasting, treating it as a cross-domain transfer learning task. These universal cross-transfer approaches, specifically recent works such as LLMTime [?] and GPT4TS [?] yield promising results in few/zero-shot forecasting approaches. These models are bootstrapped from GPT-2/3 or LLAMA-2 with suitable tokenization strategies to adapt to time-series domains.
@@
In particular, TTM is pre-trained using multiple public datasets ($\sim$ 244M samples) from the Monash data repository [^2] Accessible at https://forecastingdata.org/ [?]). Note that the datasets exhibit considerable diversity in terms of characteristics, such as the different domains, temporal resolution [^3] Resolution refers to the sampling rate of the input time series (e.g., hourly, 10 minutes, 15 minutes, etc.) (spanning from second to daily), lengths, and number of channels. Pre-training on such heterogeneous datasets cannot be handled directly by TSMixer or existing state-of-the-art (SOTA) models. Hence, TTM proposes the following enhancements to the TSMixer architecture: (i) Adaptive Patching across layers, considering the varied suitability of patch lengths for different datasets, (ii) Dataset Augmentation via Downsampling to increase coverage and samples across different resolutions, (iii) Resolution Prefix Tuning to explicitly embed resolution information in the first patch, facilitating resolution-conditioned modeling, particularly beneficial in scenarios with short history lengths. Moreover, our approach leverages multi-level modeling, where TTMs are first pre-trained in a channel-independent way and then seamlessly integrate channel mixing during fine-tuning to model target data-specific channel-correlations and exogenous infusion
@@ 2.1 Multi-level Modeling
TTM follows a multi-level architecture consisting of four key components (see Figure 1(a)): (1) The TTM Backbone is assembled using building blocks derived from the efficient TSMixer architecture [?]. TSMixer is based on simple MLP blocks that enable mixing of features within patches, across patches and channels, surpassing existing transformer-based TS approaches with minimal computational requirements. Since TSMixer is not targeted to handle multi-resolution data, we introduce various novel enhancements to it as explained later. (2) TTM Decoderfollows the same backbone architecture but is considerably smaller in size, approximately 10-20% of the size of the backbone, (3) Forecast Headconsists of a linear head designed to produce the forecast output, and (4) Optional Exogenous Mixer serves to fuse exogenous data into the model’s forecasting process. This multi-level model refactoring is required to dynamically change the working behavior of various components based on the workflow type, as explained in Section 3. In addition to the above primary components, we also have a preprocessing component as explained next.
@@ 2.2 Pre-processing
As shown in Figure 1(a) with colorless blocks, the historical time series $\bm{X}$ is first normalized per instance to have zero mean and unit standard deviation for each channel dimension, to tackle any possible distribution shifts [??]. This process is reversed at the end before computing the loss. The normalized data $\overline{\bm{X}}$ is subsequently patched $\bm{X}_{p}\in\mathbb{R}^{c\times n\times pl}$ into $n$ non-overlapping windows, each of length $pl$ and then, passed to the TTM backbone. Patching, as introduced in [?], has proven to be highly valuable for forecasting. Its effectiveness lies in preserving local semantic information, accommodating longer history, and reducing computation.
@@ 3.1 Pre-training Workflow
In the pre-training stage, we train the model on a large collection of public datasets from the Monash data repository[?]. Since the primary focus of TTM is forecasting, pre-training is modeled with a direct forecasting objective. TTM is first pre-trained in a univariate fashion with independent channels on all the existing datasets. Due to varied channel counts in pre-training datasets, modeling multivariate correlations is not feasible here; it is addressed later during fine-tuning. Multivariate pre-training datasets are initially transformed into independent univariate time series $(\bm{X}_{1},\cdots,\bm{X}_{N})\in\mathbb{R}^{c(=1)\times sl}$. These are pre-processed (Section 2.2), and subsequently fed into the TTM backbone for multi-resolution pre-training. The output of the backbone $\bm{X}_{h}^{L}\in\mathbb{R}^{(c=1)\times n\times hf}$ is passed through the decoder and forecast head to produce the forecast $\hat{\bm{Y}}\in\mathbb{R}^{(c=1)\times fl}$ which is then reverse-normalized to bring back to the original scale. We pre-train the TTM with mean squared error (MSE) loss function calculated over the forecast horizon: $\mathcal{L}=||\bm{Y}-\hat{\bm{Y}}||_{2}^{2}$. Thus for a given input context length $sl$ and forecast length $fl$, we get a pre-trained model capturing the common temporal forecasting dynamics and seasonal patterns as observed in the overall pre-training data.
Multi-Resolution Pre-training via TTM Backbone
@@
The TTM backbone consists of $L$ levels, each comprising $M$ TTM blocks with identical patch configurations. A TTM block in the $i$-th level, $i=1,\ldots,L$, receives the processed data $\bm{X}_{h}^{(i-1)}\in\mathbb{R}^{c\times n\times hf}$ from the earlier block. Each TTM block is further comprised of a patch partition model, a vanilla TSMixer block, and a patch merging block. Patch Partition block at every level $i$ increases the number of patches by a factor of $K_{i}$ and reduces the patch dimension size by the same factor by reshaping $\bm{X}_{h}^{(i-1)}\in\mathbb{R}^{c\times n\times hf}\text{to }\bm{X}_{h}^{i}\in\mathbb{R}^{c\times(n\cdot K_{i})\times(hf/K_{i})}$, where $K_{i}=2^{(L-i)}$. Note that, we set $hf=2^{m}$ for some integer $m\geq L$. Then, TSMixer is applied to the adapted data $\bm{X}_{h}^{i}$. Finally, the output from TSMixer is again reshaped to its original shape (i.e., $\mathbb{R}^{c\times n\times hf}$) in the patch merging block. Note that, as we go deeper into the network, the number of patches decreases while the patch dimension size increases leading to adaptive patching which helps in better generalization as we pre-train with multiple datasets together. This idea of adaptive patching is popular and very successful in the vision domain (E.g. Swin transformers [?] and we are the first to port it successfully to the time-series domain to resolve multi-resolution issues in pre-training with diverse TS datasets. Figure 1(b) shows the TTM backbone for $L=3$ and $M=2$. Please note that adaptive patching is enabled only in the backbone and not in the decoder, which is designed to be very lightweight.
@@
Resolution Prefix Tuning: This technique explicitly learns and incorporates a new patch embedding as a prefix into the input data based on the input resolution type (see Figure 1(b)). Similar to the concept of prefix tuning [?], this approach provides an explicit signal to the model about the resolution type for resolution-conditioned modeling. First, we map every resolution to a unique integer, which is then passed through an embedding layer to project it to the hidden dimension, $hf$. Subsequently, we expand the embedding across all channels to have a representation of shape $c\times 1\times hf$. This module is optional for the TTM backbone, particularly beneficial when the context length ($sl$) is short. In these scenarios, automatically detecting the resolution becomes a challenge for the model. Hence, by explicitly fusing the resolution information as a prefix, we can enhance the model’s ability to learn effectively across resolutions.
@@ 3.2 Fine-tuning Workflow
Zero-shot Few-shot (5%) $fl$ TTM TTM GPT4TS PatchTST TSMixer TimesNet Dlinear FEDFormer Autoformer ETTH1 96 0.365 0.366 0.543 0.557 0.554 0.892 0.547 0.593 0.681 192 0.393 0.391 0.748 0.711 0.673 0.940 0.720 0.652 0.725 336 0.415 0.421 0.754 0.816 0.678 0.945 0.984 0.731 0.761 720 0.538 - - - - - - - - ETTH2 96 0.285 0.282 0.376 0.401 0.348 0.409 0.442 0.390 0.428 192 0.341 0.338 0.418 0.452 0.419 0.483 0.617 0.457 0.496 336 0.383 0.383 0.408 0.464 0.389 0.499 1.424 0.477 0.486 720 0.441 - - - - - - - - ETTM1 96 0.413 0.359 0.386 0.399 0.361 0.606 0.332 0.628 0.726 192 0.476 0.402 0.440 0.441 0.411 0.681 0.358 0.666 0.750 336 0.553 0.424 0.485 0.499 0.467 0.786 0.402 0.807 0.851 720 0.737 0.575 0.577 0.767 0.677 0.796 0.511 0.822 0.857 ETTM2 96 0.187 0.174 0.199 0.206 0.200 0.220 0.236 0.229 0.232 192 0.261 0.240 0.256 0.264 0.265 0.311 0.306 0.394 0.291 336 0.323 0.299 0.318 0.334 0.314 0.338 0.380 0.378 0.478 720 0.436 0.407 0.46 0.454 0.410 0.509 0.674 0.523 0.553 Weather 96 0.154 0.152 0.175 0.171 0.188 0.207 0.184 0.229 0.227 192 0.203 0.198 0.227 0.230 0.234 0.272 0.228 0.265 0.278 336 0.256 0.250 0.286 0.294 0.287 0.313 0.279 0.353 0.351 720 0.329 0.326 0.366 0.384 0.365 0.400 0.364 0.391 0.387 Electricity 96 0.169 0.142 0.143 0.145 0.147 0.315 0.150 0.235 0.297 192 0.196 0.162 0.159 0.163 0.172 0.318 0.163 0.247 0.308 336 0.209 0.184 0.179 0.175 0.190 0.340 0.175 0.267 0.354 720 0.264 0.228 0.233 0.219 0.280 0.635 0.219 0.318 0.426 Traffic 96 0.518 0.401 0.419 0.404 0.408 0.854 0.427 0.670 0.795 192 0.548 0.425 0.434 0.412 0.421 0.894 0.447 0.653 0.837 336 0.55 0.437 0.449 0.439 0.477 0.853 0.478 0.707 0.867 720 0.605 - - - - - - - - TTM Zero-shot Vs SOTA Few-shot (IMP) 1% 3% 2% 31% 6% 25% 32% TTM 5% Few-shot (IMP) 12% 14% 12% 38% 17% 32% 38% Table 1: Zero-shot and Few-shot 5% performance (MSE) of TTM and all SOTA models on D1 datasets for varying forecast lengths ($fl$). Bold and underscore denote the best and the second best. For longer $fl$ (720), some datasets fail to generate fine-tuning samples with just 5% of the data, denoted as ‘-’. The last two rows show relative improvement (IMP) of TTM zero-shot and few-shot w.r.t. SOTA few-shot results. IMP is calculated based on the mean of the percentage improvement of TTM achieved across all rows.
@@ 4.1 Experimental Setting
Datasets: Pre-training employs a subset of the Monash data hub [?] of size $\sim 244$M samples. We specifically exclude datasets (like yearly, monthly) as they do not possess sufficient length for the long-term forecasting task. Moreover, we remove all the datasets that we utilize for evaluation (i.e., weather, electricity, and traffic). For zero/few-shot evaluation we consider seven public datasets (D1): ETTH1, ETTH2, ETTM1, ETTM2, Weather, Electricity & Traffic as popularly used in most prior SOTA works [??]. Since these datasets do not contain any exogenous variables nor exhibit cross-channel correlation benefits, we incorporate four other datasets (D2) for separately validating the efficacy of the decoder channel mixing and exogenous mixer module: bike sharing (BS) [?], carbon capture plant (CC) [?], and 2 more datasets from Biz-IT observability domain [?]: Application (APP) and Service (SER). Refer Appendix for full data details.
SOTA Bechmarks: We benchmark TTM with the latest public SOTA forecasting models categorized as follows: [label=()] 1. LLM-based TS pre-trained models:GPT4TS [?], LLMTime [?] 2. Self-supervised pre-trained models: SimMTM [?],Ti-MAE [?], TST [?], LaST [?], TF-C [?], CoST [?] and Ts2Vec [?] 3. TS transformer models:PatchTST [?], FEDFormer [?], Autoformer [?] 4. Other SOTA models:TSMixer [?], DLinear [?] and TimesNet [?]
Zero-shot Few-shot (10%) Data TTM (Ours) TTM (Ours) GPT4TS PatchTST TSMixer TimesNet Dlinear FEDFormer Autoformer ETTH1 0.428 0.422 0.590 0.633 0.578 0.869 0.691 0.639 0.702 ETTH2 0.362 0.360 0.397 0.415 0.370 0.479 0.605 0.466 0.488 ETTM1 0.545 0.458 0.464 0.501 0.441 0.677 0.411 0.722 0.802 ETTM2 0.302 0.278 0.293 0.296 0.282 0.320 0.316 0.463 1.342 Electricity 0.210 0.180 0.176 0.176 0.180 0.323 0.180 0.346 0.431 Traffic 0.555 0.428 0.440 0.430 0.429 0.951 0.446 0.663 0.749 Weather 0.236 0.227 0.238 0.242 0.236 0.279 0.241 0.284 0.300 TTM Zero-shot Vs SOTA Few-shot (IMP) -4% -2% -7% 27% 2% 27% 39% TTM 10% Few-shot (IMP) 7% 9% 4% 34% 13% 34% 45% Table 3: Zero-shot and Few-shot 10% performance (MSE) of TTM over all SOTA models on D1 datasets, averaged across 4 forecast lengths. The numbers with bold and underscore denote the best and the second best results respectively. The last two rows show relative improvement (IMP) of TTM zero-shot and few-shot w.r.t. SOTA few-shot results. IMP is calculated based on the mean of the percentage improvement of TTM achieved across all rows. The full table is in Appendix.